How do I calculate equilibrium national income when given? For example:
I=500 , Autonomous Consumption c= 100, marginal propensity to income b= 0.5 , G= 400 ,(X-M)= 1000
1
Expert's answer
2017-03-06T15:14:06-0500
If you need calculate equilibrium national income, we can use this formula:
Y= C+I+G+(X-M), where C= a+b*Y.
For example: we have I=500 , a= 100, b= 0.5, G= 400 ,(X-M)= 1000 that`s why
Y = 100+0,5*Y+500+400+1000
0,5Y=2000
Y=4000
ANSWER: Y=4000.